# skipabove=1em,skipbelow=1em

### v3

@@ 6.3. Social Feedback/Intervention
Social feedback/intervention refers to feedback or interventions that involve more than one person, incorporating social support or the eating environment. This type of feedback/intervention leverages group dynamics and shared experiences to reinforce closed-loop paradigms through social norms. They often incorporate interactive elements, such as family meal tracking or community network challenges, to foster collective accountability. By embedding feedback and interventions in natural social contexts, they tap into intrinsic motivations for conformity and belonging that individual ingestion-related approaches cannot address.
6.3.1. Social Feedback Only
Social feedback creates shared accountability through real-time group comparisons (e.g., displaying community members’ vegetable intake rates) or peer endorsements. It leverages social norms and observational learning—showing how similar individuals achieve balanced diets to motivate self-adjustment without prescriptive rules. This type of feedback makes healthy choices feel collectively validated rather than individually enforced. A total of 2 papers are included in this category.
@@ 6.3.1. Social Feedback Only
HealthyStadium (Eigen et al., 2018) and Crumbs (Epstein et al., 2016) utilize social feedback—through meal photo rankings and shared challenges—to encourage healthier eating habits more effectively than solo tracking. Both systems demonstrate how passive peer interaction fosters dietary mindfulness and self-adjustment. Their designs intentionally avoid direct intervention, instead relying on social dynamics such as normative comparison and gentle competition to organically shift ingestive behaviors. Both systems demonstrate the effectiveness of social feedback in promoting healthier ingestive habits through lightweight, non-intrusive engagement rather than active intervention. Critically, this lightweight, feedback-driven approach has proven to be more acceptable than traditional interventions, as users tend to resist rigid rules but embrace socially reinforced habits.
@@
Users in HealthyStadium exhibited higher engagement by striving for better health ratings, with participants consciously opting for healthier meals due to peer evaluations. One user noted, “When I was wondering whether to have a delicious meal or a healthy meal, I chose a healthy one” (Eigen et al., 2018). Meanwhile, Crumbs leveraged daily challenges and social interaction, with 72% of users enjoying the challenges and 46% expressing intent to continue using the platform. Additionally, the social group reported motivational effects, such as ”not wanting to fall behind others” (Epstein et al., 2016).
6.3.2. Social Intervention Only
@@
Social intervention for healthy ingestion actively leverages interpersonal dynamics to reshape dietary behaviors through actual engagement. This type of intervention employs real-time social triggers—such as parent-child tactile feedback—that directly mediate ingestive actions within relational contexts. By embedding behavioral guidance within social interactions (e.g., family meals and shared environments), these interventions transform abstract nutritional awareness into contextually anchored behavior change. A total of 4 papers are included in this category.
@@
The interactive tableware system (Joi et al., 2016) successfully encouraged vegetable consumption in children aged 3–4, with observed behavioral shifts, such as a 4-year-old transitioning from being spoon-fed to self-feeding. Additionally, it improved parent-child interaction through gamified learning. Similarly, a middle-school healthy eating intervention demonstrated gender-dependent effects, notably reducing fat intake in girls—particularly when parents were involved—though it had a limited impact on fruit and water intake as well as soft drink consumption (Haerens et al., 2007). Meanwhile, the auto-generated postcard system facilitated self-reflection, leading to expressed intent to modify behaviors (Sun et al., 2020).
@@
6.3.3. Social Feedback with Intervention
Social feedback with intervention uniquely hybridizes data-driven insights with social dynamics to synergistically promote healthy ingestive behaviors. These approaches integrate real-time nutritional tracking and visualization with socially embedded prompts—such as dietary monitoring applications for parents (Alloghani et al., 2016) or office hydration competitions (Chiu et al., 2009)—creating a dual mechanism for behavior change. This ensures that targeted nudges are both empirically grounded and socially reinforced. A total of 2 papers are included in this category.
Regarding the outcomes of this approach, the mobile health monitoring application successfully enhanced user trust and acceptance by combining QR-code food tracking, geo-location, and wearable sensors. Users prioritized ease of use and security as key determinants of perceived usefulness (Alloghani et al., 2016). Similarly, Playful Bottle leveraged social reminders alongside system-based hydration games, demonstrating that social prompts were more effective than automated reminders. This approach not only increased water intake volume but also improved regularity and response speed among office workers (Chiu et al., 2009). Notably, social elements consistently outperformed system-only interventions. Users of Playful Bottle responded faster and more reliably to peer reminders than to impersonal alerts (Chiu et al., 2009), echoing findings from the obesity app study that indicated trust and usability are heightened by socially embedded design choices (Alloghani et al., 2016).
6.4. Other Specific Feedback/Intervention
Other specific approaches not covered above are categorized into three key themes: Perceptual Augmentation (i.e., enhancing sensory engagement with food), Mindful and Psychological (i.e., fostering awareness and emotional connections to ingestion), and Smart Monitoring (i.e., automated tracking systems). These frameworks illustrate a range of technological strategies for promoting changes in ingestive behavior.
@@ 6.4.1. Perceptual Augmentation
Perceptual augmentation actively modifies ingestion experiences by technologically enhancing sensory feedback during consumption. These approaches employ real-time cross-modal stimulation—such as altering chewing sounds to influence flavor perception (Kleinberger et al., 2023) or using augmented reality (AR) to visually enhance food texture and taste (Koizumi et al., 2011; Narumi et al., 2011)—to directly reshape how food is sensorially experienced. By dynamically manipulating auditory, visual, or haptic cues during the act of ingestion, such interventions create immersive sensory modifications rather than relying on traditional feedback or interventions. A total of 3 papers are included in this category.
Emerging perceptual augmentation technologies demonstrate how non-invasive sensory enhancement can effectively alter ingestive behaviors and perceptions. Auditory feedback systems leverage cross-modal effects to enhance flavor perception—significantly improving texture ratings such as crispiness and reducing hunger (Kleinberger et al., 2023; Koizumi et al., 2011). Meanwhile, Metacookie+ utilizes AR-driven visual and olfactory cues to manipulate taste perception, achieving a 79.3% success rate in altering familiar flavors, although unfamiliar tastes, such as mushroom, saw lower recognition at 42.9% (Narumi et al., 2011).
@@ 6.4.2. Mindful & Psychological
Mindful and psychological approaches reshape ingestive behaviors by targeting cognitive-emotional processes and fostering real-time awareness during consumption. These interventions employ evidence-based techniques, such as future-self visualization to motivate long-term behavior change (Rho et al., 2017), emotion-aware wearables to detect and interrupt emotional eating (Carroll et al., 2013), and conversational agents to guide mindful food preparation (Parra et al., 2023). These strategies directly address the psychological roots of maladaptive eating patterns. Ultimately, these interventions cultivate mindful eating by transforming abstract nutritional knowledge into embodied psychological awareness—where every bite becomes an opportunity to nourish both body and mind, and where present-moment consciousness bridges our current selves with the healthier future selves we aspire to become. A total of 4 papers are included in this category.
Mindful and psychological approaches demonstrate strong potential in promoting healthier ingestive behaviors. FutureSelf effectively motivated users by forecasting weight outcomes, with participants reporting increased goal awareness and behavioral conscientiousness (Rho et al., 2017). EmoTree’s Just-In-Time interventions improved emotional eating awareness by 37.5% and received an averaged rating of 3 out of 5 for effectiveness in stress reduction, although its impact on preventing eating was limited, with an averaged rating of 2.2 out of 5 (Carroll et al., 2013). These interventions succeed by leveraging introspection and emotional engagement but require refinement in adaptive personalization and long-term habit formation (Carroll et al., 2013; Rho et al., 2017). Future work should aim to balance immediate support with sustained behavioral reinforcement.
@@ 6.4.3. Smart Monitoring
Smart monitoring leverages wearable technologies and ambient sensors—from alcohol-detecting ice cubes (Matsui et al., 2018) to eyeglass-mounted detectors (Bedri et al., 2022) and wrist-worn motion trackers (Ye et al., 2016)—to autonomously capture ingestive behaviors (e.g., drinking and chewing) and environmental interactions (e.g., food selection (Volkova et al., 2016)). This approach transforms passive observation into actionable behavioral datasets. A total of 4 papers are included in this category.
Key strengths in this category include reduced cognitive load (via automated detection) and enhanced self-awareness, particularly regarding snack consumption (Bedri et al., 2022; Ye et al., 2016). However, challenges remain, such as false positives—evidenced by 50% of users receiving 6 or more unnecessary alerts in FitNibble (Bedri et al., 2022)—and issues related to device wearability. Smart monitoring systems have demonstrated effectiveness in improving dietary tracking compliance. FitNibble reduced missed logs by 19.6% and lowered journaling difficulty (Bedri et al., 2022). Similarly, the wrist-worn Pebble system received positive usability scores (2.7 on a scale from -3 to 3), particularly benefiting irregular eaters through visual logging prompts (Ye et al., 2016).
@@ 7. Discussion
We synthesize insights from the reviewed systems to reflect on broader limitations and opportunities in the design of behavioral sensing and feedback/intervention in ingestion health. By revisiting the conceptual and technical landscape through the lens of system completeness, behavioral scope, and interdisciplinary integration, we offer a summary analysis that responds to our fourth research question. The discussion is organized around three key challenges: (1) the difficulty of accurately measuring and intervening in real-world ingestive behavior, (2) the disconnect between sensing-focused and intervention-driven research efforts, and (3) the methodological and epistemological gap between HCI and clinical health domains. For each, we outline design directions aimed at informing future development of more effective and ecologically embedded closed-loop systems.
@@ 7.1. Challenges in Measuring and Intervening in Real-World Ingestive Behavior
Despite recent advances in sensing technologies and intervention strategies, significant challenges remain in accurately measuring and applying ingestion behavior interventions in real-world settings. These reflections are informed by our previous analysis of intervention effectiveness and user acceptance (Chapter  6), where we observed overall evaluation methods and user feedback in ecological validity, evaluation consistency, and behavioral coverage. In particular, the absence of robust and context-sensitive evaluation frameworks continues to hinder the translation of promising closed-loop sensing–intervention systems into everyday practice. These challenges can be broadly categorized into three areas:
@@ 7.1.1. Validity and Consistency of Ingestive Behavior Measurement Indicators
As observed in our analysis of feedback and intervention effectiveness (Chapter 6), the lack of consistency in how ingestive behaviors are defined and evaluated poses a major barrier to assessing and comparing intervention outcomes. Although many studies focus on similar behavioral constructs, the measurement indicators they employ vary widely, resulting in inconsistencies across interpretation, validation, and generalization. Traditional instruments including questionnaires and scales are prevalent in the field of ingestion health intervention, such as the Three-Factor Eating Questionnaire (TFEQ) (Stunkard and Messick, 1985), which assesses cognitive restraint, disinhibition and hunger, and the Eating Behavior Inventory (EBI) (O’Neil et al., 1979), which targets weight-control behaviors. Despite their widespread use, their applicability across diverse populations and contexts has still been questioned, remaining non-universal; for emerging eating behavior interventions that rely on sensing and computational methods, the challenge is further compounded by diverse evaluation metrics and the lack of standardized assessment frameworks. Even when interventions address the same behavioral target—such as eating speed or intake pace—the metrics used to quantify outcomes are highly heterogeneous. To date, medical literature offers no standardized clinical definition for many of these behavioral constructs. For example, although rapid eating is commonly linked to adverse health outcomes, there is no standardized definition or measurement approach; current systems rely on varied metrics such as chewing rate, swallowing frequency, bite intervals, or total meal duration to approximate eating speed. This diversity highlights not only methodological flexibility but also the absence of a shared ground truth in computationally mediated behavior evaluation. Developing valid, standardized, and context-aware behavioral metrics is therefore a critical direction for ensuring scientific rigor and comparability in ingestion health research.
@@
A substantial portion of ingestive behavior research has been conducted in controlled laboratory environments, which, though offering experimental rigor, often fail to reflect the complexity of real-world contexts. Lab protocols typically exclude environmental, social, and emotional factors that influence natural eating, leading to potential over- or underestimation of intervention effectiveness when deployed outside the lab (De Castro, 2000; Bell et al., 2020; Hiraguchi et al., 2023).
For instance, bite and chewing detection systems are often validated in constrained conditions—such as fixed posture, single food types, and minimized distraction—within laboratory settings [e.g., (Scisco et al., 2011; Kim et al., 2016b; Hori et al., 2023; Zhang et al., 2019)]. However, in real-world environments, behaviors are shaped by multitasking, social interaction, and contextual variability, which can significantly impact system performance and user adherence (Bell et al., 2020; Hiraguchi et al., 2023; Jansen et al., 2022). Moreover, some systems using head-mounted cameras, chest-worn sensors, or bulky wearables often perform well in short-term lab trials, but can feel intrusive or uncomfortable in daily life, potentially disrupting natural behavior and lowering compliance (Bedri et al., 2022; Bell et al., 2020). These real-world deployment issues, such as behavioral reactivity and long-term disengagement, are frequently underestimated in lab-based evaluations (Jansen et al., 2022). Bridging this gap calls for validation in situation that accounts for behavioral variability, contextual complexity, and sustained user acceptance (Hiraguchi et al., 2023).
7.1.3. Potential Adverse Effects of Intervention Deployments
While designed to promote healthier behaviors, eating interventions may produce unintended negative effects. A key concern is ”food noise”, which means persistent and intrusive thoughts about eating triggered by constant monitoring. It can elevate anxiety and contribute to disordered eating patterns (Hayashi et al., 2023). Meanwhile, interventions requiring user interaction during meals, such as smart utensils or app prompts, may also disrupt social interactions or conflict with cultural norms, especially in shared settings (Robinson et al., 2020; Kadomura et al., 2014). Moreover, long-term use of intrusive devices like head-mounted sensors or chewing monitors has been linked to discomfort and user dropout (Bedri et al., 2022). These risks underscore the need for balancing behavioral influence with psychological comfort and contextual fit.
@@ 7.2. Gaps between Sensing-Based and Intervention-Based Studies on Ingestive Behavior
Recent reviews on sensor-based research in ingestive behavior [e.g., (Hiraguchi et al., 2023; Bell et al., 2020)] have highlighted that while the target behaviors addressed in sensing studies overlap with those identified in intervention-focused research, significant gaps remain between the two approaches. These gaps primarily stem from differences in modality alignment and the application of behavior change strategies.

### v4

@@ 6.3. Social Intervention
Social intervention refers to behavioral strategies that leverage group dynamics and social contexts (e.g., family meals or community networks) to drive change and reinforce the closed-loop paradigms. By incorporating shared experiences and social norms, this approach fosters collective accountability and taps into intrinsic motivations for belonging — aspects that individual-focused interventions cannot address.
6.3.1. Passive Social Intervention
Passive social intervention fosters accountability through real-time peer comparisons, leveraging social norms and observational learning to motivate self-regulation while making healthy choices feel socially validated. A total of 2 papers are included in this category.
@@ 6.3.1. Passive Social Intervention
HealthyStadium (Eigen et al., 2018) and Crumbs (Epstein et al., 2016) utilize passive social intervention — through meal photo rankings and shared challenges — to encourage healthier eating habits more effectively than solo tracking. Both systems validate that passive social intervention cultivates dietary self-regulation through non-intrusive peer engagement, rather than active interventions. Critically, this lightweight, feedback-driven approach has proven to be more acceptable than traditional interventions, as users tend to embrace socially reinforced habits.
@@
Users in HealthyStadium exhibited higher engagement by striving for better health ratings, with participants consciously opting for healthier meals due to peer evaluations (Eigen et al., 2018). Meanwhile, Crumbs leveraged daily challenges and social interaction, with 72% of users enjoying the challenges and 46% expressing intent to continue using the platform. Moreover, the social group reported motivational effects, such as ”not wanting to fall behind others(Epstein et al., 2016).
6.3.2. Active Social Intervention
@@
Active social intervention for healthy ingestion leverages interpersonal dynamics to reshape dietary behaviors through actual engagement. This type of intervention employs real-time social triggers that directly mediate ingestive actions within relational contexts. By embedding behavioral guidance within social interactions (e.g., family meals and shared environments), these interventions transform abstract nutritional awareness into contextually anchored behavior change. A total of 4 papers are included in this category.
@@
The interactive tableware system (Joi et al., 2016) increased vegetable consumption in 3-4-year-olds, demonstrating behavioral shifts like transitioning from spoon-feeding to self-feeding while enhancing parent-child interactions through gamification. Similarly, a school-based intervention (Haerens et al., 2007) showed gender-specific effects, significantly reducing girls’ fat intake. Complementing these, the automated postcard system (Sun et al., 2020) also effectively promoted self-reflection and behavior change intentions.
@@
6.3.3. Combined Social Intervention
Combined social intervention uniquely hybridizes data-driven insights with social dynamics to synergistically promote healthy ingestive behaviors. These approaches integrate real-time nutritional tracking and visualization with socially embedded prompts — such as dietary monitoring applications for parents (Alloghani et al., 2016) or office hydration competitions (Chiu et al., 2009) — creating a dual mechanism for behavior change. This ensures that targeted nudges are both empirically grounded and socially reinforced. A total of 2 papers are included in this category.
Regarding the outcomes of this approach, the mobile health app (Alloghani et al., 2016) boosted user trust and acceptance by integrating QR-code tracking, geo-location, and wearable sensors, with ease of use and security being key determinants of perceived usefulness. Similarly, Playful Bottle (Chiu et al., 2009) demonstrated that social reminders were more effective than automated alerts, improving both the volume and regularity of water intake among office workers. These studies collectively show that socially-embedded interventions consistently outperform system-only approaches, with peer-based prompts generating faster responses and socially-aware designs enhancing both usability and trust (Chiu et al., 2009; Alloghani et al., 2016).
6.4. Other Specific Interventions
Other specific approaches not covered above are categorized into three key themes: Perceptual Augmentation, Mindful and Psychological, and Smart Monitoring. These frameworks illustrate a range of technological strategies for promoting changes in ingestive behavior.
@@ 6.4.1. Perceptual Augmentation
Perceptual augmentation technologies modify ingestion experiences through real-time cross-modal stimulation, such as altering chewing sounds to enhance flavor perception (Kleinberger et al., 2023) or using AR to visually manipulate texture and taste (Koizumi et al., 2011; Narumi et al., 2011). These non-invasive sensory enhancements effectively alter ingestive behaviors—auditory feedback improves texture ratings and reduces hunger, while AR-driven cues (e.g., Metacookie+) achieve a 79.3% success rate in modifying familiar flavors, albeit with lower recognition for unfamiliar tastes (42.9%) (Narumi et al., 2011). A total of 3 papers are included in this category.
@@ 6.4.2. Mindful & Psychological
Mindful and psychological approaches reshape ingestive behaviors by targeting cognitive-emotional processes through techniques like future-self visualization (Rho et al., 2017), emotion-aware wearables (Carroll et al., 2013), and conversational agents (Parra et al., 2023). These interventions foster real-time awareness, with FutureSelf increasing goal consciousness and EmoTree improving emotional eating awareness by 37.5%. While effective in promoting introspection and short-term behavioral change, they require refinement in adaptive personalization and long-term habit formation (Carroll et al., 2013; Rho et al., 2017). A total of 4 papers are included in this category.
@@ 6.4.3. Smart Monitoring
Smart monitoring utilizes wearable and ambient sensors — including alcohol-detecting ice cubes (Matsui et al., 2018), eyeglass-mounted detectors (Bedri et al., 2022), and wrist-worn trackers (Ye et al., 2016) — to autonomously capture ingestive behaviors and environmental interactions (Volkova et al., 2016). While demonstrating effectiveness (FitNibble reduced missed logs by 19.6% and improved journaling ease (Bedri et al., 2022), while Pebble scored 2.7 on a scale from -3 to 3 for usability (Ye et al., 2016)), challenges persist (50% of FitNibble users received more than 6 false alerts (Bedri et al., 2022)), highlighting wearability and accuracy tradeoffs. A total of 4 papers are included in this category.
@@ 7. Discussion
We synthesize insights from the reviewed systems to identify challenges and opportunities in designing a closed-loop paradigm for ingestion health, drawing on system completeness, behavioral scope, and interdisciplinary integration to address our fourth research question. The discussion centers on three key challenges: (1) the difficulty of accurately measuring and intervening in real-world ingestive behavior, (2) the disconnect between sensing technologies and intervention practices, and (3) the methodological and epistemological gap between HCI and clinical health domains. For each, we summarize the underlying issues and propose design principles to support the development of more effective, ecologically embedded closed-loop systems.
@@ 7.1. Challenges in Measuring and Intervening in Real-World Ingestive Behavior
Despite rapid advances in sensing technologies and intervention strategies, accurately measuring and intervening in real-world ingestion behavior remains challenging. Our analysis of intervention effectiveness and user acceptance in Chapter  6 revealed the absence of ecological validity, consistency, and behavioral coverage. The lack of robust and context-sensitive evaluation frameworks makes it difficult for promising closed-loop systems to integrate into daily life. This challenge can be further unpacked into the following aspects.
@@ 7.1.1. Validity and Consistency of Ingestive Behavior Measurement Indicators
Our analysis of intervention effectiveness revealed that studies frequently apply inconsistent operational definitions and behavioral metrics, even when addressing the same ingestive behavior construct-hindering interpretation, validation, and cross-study comparability. In traditional research of ingestion health, validated questionnaire tools such as the Three-Factor Eating Questionnaire (TFEQ) (Stunkard and Messick, 1985) and the Eating Behavior Inventory (EBI) (O’Neil et al., 1979) are commonly used to assess psychological and behavioral traits. While these instruments are well established, they are not well aligned with emerging context-aware approaches. Studies using computational methods often adopt their own evaluation criteria, lacking a unified ”gold standard”. For instance, in measuring a common behavioral target such as eating speed, no standardized definition currently exists. Current systems approximate eating speed using a range of metrics, such as chewing rate, swallowing frequency, bite intervals, or total meal duration. This diversity, while reflecting methodological flexibility, also underscores a lack of consistency and validation, which limits cross-study comparability and introduces potential bias in evaluating intervention outcomes. Developing valid, standardized, and context-sensitive behavioral indicators is therefore essential to ensuring scientific rigor and cross-study comparability in ingestion health research.
@@
A substantial portion of researches on ingestion health is still conducted in controlled laboratory settings. While offering rigor, these settings simplify reality by excluding environmental, social, and emotional factors-facilitating statistical detection but risking over- or underestimation of effectiveness when interventions are applied in real-world contexts (De Castro, 2000; Bell et al., 2020; Hiraguchi et al., 2023). Bite and chewing detection systems, for instance, are typically validated under constrained lab conditions — fixed posture, limited food types, and minimal distraction [e.g., (Scisco et al., 2011; Kim et al., 2016b; Hori et al., 2023; Zhang et al., 2019)]. However, in real-world settings, multitasking, social interaction, and context variability can significantly affect system performance and adherence (Bell et al., 2020; Hiraguchi et al., 2023; Jansen et al., 2022). Wearable systems, such as head-mounted cameras or chest sensors, may perform well in short lab trials, but often feel intrusive in daily life, potentially disrupting natural behavior and lowering compliance (Bedri et al., 2022; Bell et al., 2020). These real-world deployment issues, such as behavioral reactivity and long-term disengagement, are frequently underestimated in lab-based evaluations (Jansen et al., 2022). Bridging this gap calls for validation in situation that accounts for behavioral variability, contextual complexity, and sustained user acceptance (Hiraguchi et al., 2023).
7.1.3. Potential Adverse Effects and Privacy/Security Concerns of Intervention Deployments
While designed to promote healthier behaviors, ingestion interventions may lead to unintended side effects. A key concern is ”food noise” — persistent eating-related thoughts triggered by continuous monitoring, which may increase anxiety and contribute to disordered patterns (Hayashi et al., 2023). Meanwhile, interventions involving in-meal interaction (e.g., smart utensils, app prompts) may disrupt social dynamics or clash with cultural norms, especially during shared meals (Robinson et al., 2020; Kadomura et al., 2014). Moreover, intrusive devices (e.g., head-mounted sensors, chewing monitors) may cause discomfort and user dropout over time (Bedri et al., 2022). These risks underscore the need to balance behavioral influence with psychological comfort and contextual fit.
Finally, privacy and security pose critical ethical concerns for real-world intervention deployment. Continuous sensing systems collect highly sensitive health and behavioral data, raising ethical concerns over encryption, anonymization, and transparency in data use (Radanliev, 2025). However, with little access to concrete context-aware computing methods in the background, users often struggle to establish trust during usability evaluations (Faltaous et al., 2021). Various data-collection modalities with tough methods may provoke user discomfort or resistance (Islam and Bae, 2024; Gao et al., 2023; Mollyn et al., 2022). Thus, ensuring data privacy and security is not only ethically imperative but essential for sustained user engagement. Together, these concerns call for a user-centered approach to intervention design — one that maximizes behavioral impact without compromising user comfort, privacy, or dignity.
@@ 7.2. Gaps between Sensing-Based and Intervention-Based Studies on Ingestive Behavior
Recent reviews on sensor-based research in ingestive behavior [e.g., (Hiraguchi et al., 2023; Bell et al., 2020)] reveal a clear gap between sensing-driven and intervention-oriented studies. While both address overlapping behavioral targets, they differ significantly in modality integration and the use of behavior change strategies.